OpenStudy (anonymous):

A spinner and a fair number cube are used in a bet. The spinner has an equal chance of landing on 1 of 4 colors purple, green, blueborbyellow,the faces of the cube are labeled 1 through 6. What is the probability of a player spinning the color green and then rolling a 5 or 6

OpenStudy (kropot72):

P(green) = 1/4 P(5 or 6) = 2/6 P(green + 5 or 6)= 1/4 * 2/6